OpenStudy (anonymous):

phase shift? y=-2cos(4x+pi)

OpenStudy (sidsiddhartha):

take a standard form y=Acos(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B

OpenStudy (sidsiddhartha):

just compare with the function given

OpenStudy (anonymous):

pi/4?

OpenStudy (sidsiddhartha):

yes to the left

OpenStudy (anonymous):

thanks

OpenStudy (sidsiddhartha):

:)